OpenStudy (anonymous):

Solve -3x2 - 4x - 4 = 0

OpenStudy (anonymous):

I got x=2+-2isqr2/3 is this correct?

OpenStudy (anonymous):

HELP?!

OpenStudy (chmvijay):

its -3x2 or -x2

OpenStudy (anonymous):

its -3^2

OpenStudy (anonymous):

and these are the answer choices: x equals quantity of 2 plus or minus 4i square root of 2 all over 3 x equals quantity of 2 plus or minus 2i square root of 2 all over 3 x equals quantity of negative 2 plus or minus 2i square root of 2 all over 3 x equals quantity of negative 2 plus or minus 4i square root of 2 all over 3

OpenStudy (anonymous):

-3x2-4x-4=0 -6-4=0+4x -10=4x x=-10/4